OpenStudy (anonymous):

11. A racing car accelerates uniformly through three gears, changes with the following average speed: 20 ms-1 for 2.0 s 40 ms-1 for 2.0 s 60 ms-1 for 6.0 s What is the overall average speed of the car ?

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

Ya ur answer is correct.

OpenStudy (anonymous):

ty :) I need ur help in one more problem 13. A generator produces 100 kW of power at a potential difference of 10 kV. The power is transmitted through cables of total resistance 5 ohm . How much power is dissipated in the cables?

OpenStudy (anonymous):

@oksuz_ 13 question ?????

OpenStudy (anonymous):

Power dissipated formula is P = I^2R ...which formula I use for finding 'I ' value ????

OpenStudy (nurali):

If the generator is delivering 100 KW of power and the Voltage is 10 KV, then it's supplying current at I = P/E = (100,000/10,000) = 10 Amperes. If 10A is flowing through cable with a total resistance of 5 ohms, the power dissipated by the cable is P = I^2R = (10)^2 x (5) = 500 watts.

OpenStudy (anonymous):

ty @Nurali :D

OpenStudy (nurali):

Anytime.

OpenStudy (gtxmuqsit):

the answer for above qts is 48 m/s2